Date: Tue, 7 May 2002 11:43:31 GMT From: jurek@apo.nmsu.edu (Jurek Krzesinski) Subject: 2.5m obslog 20020506 (MJD 52401) APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday May 6, 2002 (MJD 52401) ---=== OBSERVING TEAM ===--- swing: Dan Long night: Atsuko Nitta Kleinman Jurek Krzesinski helping around: Scot Kleinman ---=== OBSERVING PLAN ===--- Science! Spectroscopy... ---=== OBSERVING SUMMARY ===--- Most of the night it was cloudy, but we managed to get 2 plates (956 & 796) done. Also we got some exposures on other two plates done. Nothing in the problem section tonight. ---=== OBSERVING LOG ===--- Afternoon checkout: ok. Clouds ... - but maybe we can get some photons tonight. 01:40Z accidently killed xdm on host... - Craig helped us to solve the problem. 02:10Z Telescope opened, cart 1 plt 796 mounted. 02:53Z fk5 star on slew. Clouds... plate NOT done. 06:31Z cart 2 plt 956 - mounted, stars on slew. The strange thing happened: there was guider fiber number 1 shown on the top of fiber 10-th window in the saaoimage... We also spot a big bad pixel spot in the fiber 10-th flat field image, on that cart - could someone check it, please? Plate done with three exposures, but all smears were yellow and red. 07:55Z cart 4 plt 918 mouted, stars on slew. Clouds on the first exposure... NOT done. 10:35Z cart 1 plt 796 mounted, stars on slew. Trying to get an exposure, before sun rise ;> ... DONE !!! :))) endNight - ok. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields nGood rowMean rowSig colMean colSig rot --------------------------------------------------------------- ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 00:27 14929 0.0 bias 00:54 14930 0.0 bias ----- sequence 14931, plate 956 ------- 06:37 14931 10.0 flat 06:39 14932 2.0 arc 07:12 14933 1202.0 target 07:30 14934 901.0 target 07:43 14935 601.0 target 07:49 14936 240.0 smear ----- sequence 14937, plate 918 ------- 08:02 14937 10.0 flat 08:04 14938 2.0 arc 08:25 14939 901.0 target 08:50 14940 1055.0 target ----- sequence 14941, plate 796 ------- 10:43 14941 10.0 flat 10:45 14942 2.0 arc 11:03 14943 900.0 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 06:40Z 2 956 206.89 0.0008 56.41 0.0006 49.48 -0.0065 1.000160 08:10Z 4 918 -38.67 0.0010 48.02 0.0001 148.22 -0.0244 0.999990 10:51Z 1 796 227.97 0.0007 51.90 0.0016 80.07 -0.0040 1.000210 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL5298 Stays: JL5299 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 07:05Z 2 956 1.00016 -1579 -339 -200 210 53.9 10.5 10 270 gdr 1.6 08:19Z 4 918 0.99999 99 1123 -150 -42 46.7 10.3 8 276 gdr 1.5 10:50Z 1 796 1.00021 -2073 -718 -175 228 52.0 9.8 14 243 gdr 1.8 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:23Z 59 14 17 264 (W) 1729 - 00:55Z 56 11 13 266 (W) 1513 - 01:26Z 53 8 16 262 (W) 1681 - 02:35Z 52 7 10 266 (W) 1860 - 03:05Z 53 8 12 268 (W) 1759 - 03:35Z 53 8 13 270 (W) 1602 - 04:06Z 52 6 13 274 (W) 1594 - 04:36Z 52 7 10 267 (W) 1422 - 05:06Z 52 7 16 265 (W) 1437 - 05:36Z 53 7 7 272 (W) 1429 - 06:08Z 52 7 14 269 (W) 1492 - 06:39Z 52 6 9 268 (W) 1498 - 07:09Z 51 5 6 266 (W) 1519 - 07:39Z 50 5 11 261 (W) 1565 - 08:10Z 50 5 9 266 (W) 1564 - 08:42Z 50 5 11 260 (W) 1518 - 09:14Z 49 5 12 260 (W) 1488 - 09:46Z 47 2 13 264 (W) 1648 - 10:19Z 47 2 14 246 (WSW) 1579 - 10:49Z 49 3 14 252 (WSW) 1581 - 11:19Z 48 2 10 251 (WSW) 1688 - ---=== TELESCOPE STATUS ===--- 02:10Z Telescope opened, dewars refilled. ---=== SOFTWARE USED ===--- IOP/SOP: v3_78_0 Watcher: v2_17_0 MCP: v5_11_3 TPM: v2_8_0 AstroDa: v14_47 TCC: 2.6.2 sdssProcedures: v1_44 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.0770 0.0760 Axial B 0.7970 0.7990 Axial C 0.0720 0.0810 Trans D -1.5160 -1.5180 Lateral E 1.8360 1.8240 Lateral F 1.4040 1.3920 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES PriDesOrient: 0.00 -11.80 23.00 1256.90 642.10 PriOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.00 Air Temp.: 12.7 Alt.: 29.998725 MIGS TONIGHT NOMINAL Axial A 1.5490 1.5500 Axial B 1.2200 1.2220 Axial C 1.2140 1.2140 Trans D -0.5750 -0.5770 GALILS Commanded: 1599567. 1565293. 1586288. -10400. -9750. Actual: 1599600. 1565250. 1586350. -10400. -9750. SETMIR VALUES SecDesOrient: 1257.00 -3.40 -10.00 0.00 133.82 SecOrient: 1256.98 -3.38 -9.98 -0.22 133.68 ---=== PROBLEMS IN DETAIL ===---